OpenStudy (akohl):

@Abhisar Planet Z has a rocky surface. At height 3240 km above the surface, the gravitational acceleration is 48 percent of its value at the surface What is the radius of the planet? 7310 km 9120 km 8280 km 6530 km

OpenStudy (akohl):

@aaronq

OpenStudy (anonymous):

Let the radius of the planet be R and mass of the planet be M, then gravitational acceleration at a distance d from the surface is \[g = G\frac{M}{(R+d)^2}\] g at the surface i.e at d = 0 is \[g_0 = G\frac{M}{R^2}\] Since \(g\) is 48% of \(g_0\) at d = 3240 km we have \(g = \frac{48}{100} \times g_0\) \[G\frac{M}{(R+3240)^2} = \frac{48}{100} \times G\frac{M}{R^2}\] Can you solve this for R ?

OpenStudy (akohl):

There's no m though

OpenStudy (akohl):

Also what's G value

OpenStudy (akohl):

@Abhisar can u help finish this

OpenStudy (anonymous):

You don't need G and as they will cancel out !

OpenStudy (anonymous):

You have \((R+3240)^2 = \frac{100}{48} \times R^2\) Can you solve for R now ?

OpenStudy (akohl):

Should it b 48/100 not the reciprocal

OpenStudy (anonymous):

It should be 100/48

OpenStudy (anonymous):

Can you do the simplification further and solve for R?

OpenStudy (akohl):

I solved and got 3113

OpenStudy (akohl):

That's not an answer though

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

It comes out to be \(R = \frac{6400}{13} \times (6 + 5\sqrt3)\) which is around 7217 km. So you can choose first option.

OpenStudy (anonymous):

Sorry, *6480, Now you get 7307 km

OpenStudy (akohl):

Thanks